OpenStudy (shaeelynn):

The number of buckets of paint n needed to paint a fence varies directly with the total area a of the fence and inversely with the amount of paint p in a bucket. It takes three 1-gallon buckets of paint to paint 72 square feet of fence. How many 1-gallon buckets will be needed to paint 90 square feet of fence?

OpenStudy (robtobey2):

Ceiling[3*(90/72)] = 4

OpenStudy (shaeelynn):

is that the answer?

OpenStudy (robtobey2):

4 is the answer. It is assumed that you have to buy an integral number of paint cans. Ceiling is a function from Mathematica that "gives the smallest integer greater than or equal to x."

OpenStudy (robtobey2):

\[\frac{90\ 3}{72}=\frac{15}{4}=3+\frac{3}{4}=3.75=4 \text{ rounded} \text{ up} \]